OpenStudy (anonymous):

A race horse took 2 2/5 minutes to run 1 1/2 miles. How long would it take the horse to run one mile? A. 9/8 minute B. 1 1/3 minutes C. 1 1/10 minutes D. 1 2/5 minutes

OpenStudy (texaschic101):

2 2/5 = 12/5 1 1/2 = 3/2 (12/5) / (3/2) = x/1 cross multiply 3/2x = 12/5 -- divide by 3/2 x = (12/5) / (3/2) x = 12/5 * 2/3 x = 24/15 x = 1 9/15 which reduces to 1 3/5 minutes hmmm...I am getting something different